# Constrained Black-Box Attacks Against Multi-Agent Reinforcement Learning

### v1

@@ Abstract
Collaborative multi-agent reinforcement learning (c-MARL) has rapidly evolved, offering state-of-the-art algorithms for real-world applications, including sensitive domains. However, a key challenge to its widespread adoption is the lack of a thorough investigation into its vulnerabilities to adversarial attacks. Existing work predominantly focuses on training-time attacks or unrealistic scenarios, such as access to policy weights or the ability to train surrogate policies. In this paper, we investigate new vulnerabilities under more realistic and constrained conditions, assuming an adversary can only collect and perturb the observations of deployed agents. We also consider scenarios where the adversary has no access at all. We propose simple yet highly effective algorithms for generating adversarial perturbations designed to misalign how victim agents perceive their environment. Our approach is empirically validated on three benchmarks and 22 environments, demonstrating its effectiveness across diverse algorithms and environments. Furthermore, we show that our algorithm is sample-efficient, requiring only 1,000 samples compared to the millions needed by previous methods.
Index Terms:
Multi-agent reinforcement learning, adversarial attack, black box attack
I Introduction
@@
Collaborative multi-agent reinforcement learning (c-MARL) algorithms have demonstrated state-of-the-art performances in complex cooperative tasks [26, 29], making them well-suited for solving real-world problems across various domains [14, 21, 32]. However, a critical prerequisite for the widespread adoption of c-MARL is a full understanding of its vulnerabilities to adversarial attacks [7, 9], particularly when deployed.
@@
While much of the literature on adversarial attacks against c-MARL focuses on training-time attacks [33, 13, 2, 6], we focus instead on test-time attacks, where the adversary is present during deployment. Prior work on test-time attacks [23, 11, 17] has primarily considered white-box threat models, in which the adversary has access to the policy architecture and its parameters. This scenario is not always feasible. Moreover, we argue that having such access can be considered a successful attack in itself, as it typically represents proprietary knowledge with significant financial implications if leaked. In contrast, black-box threat models [7] do not assume access to the policy’s weights or its architecture; instead, they often involve learning a surrogate policy network. This can be achieved either by training the policy from scratch [7] or through imitation learning [28, 8]. The former requires access to the training environment, while the latter relies on having access to both observations and actions (i.e., expert demonstrations) or the ability to query the model. But what if the adversary is prevented from having such access to train the surrogate policy?
@@
In our work, we examine test-time attacks against c-MARL in black-box settings, but we push the standard assumptions a step further: we consider an adversary whose access is restricted to the observations (no access to the actions) of deployed agents, referred to as infected agents. The adversary does not have full control over these infected agents; instead, they can only add small perturbations to their observations. To the best of our knowledge, we are the first to investigate adversarial attacks under such constrained conditions.
Figure 1: Assumption Strength vs Plausibility: There is an inverse relationship between the strength of a threat model’s assumptions and its practical relevance: the less access an attacker is assumed to have, the more plausible the scenario becomes in real-world settings.
The first main contribution of our paper is to answer the question of whether an adversary with such limited information and capabilities can sabotage a c-MARL system and, if so, how they might achieve this. Not only is the answer yes, but the adversary can inflict significant damage with as few as 1,000 collected samples, in stark contrast to previous approaches requiring millions of samples [23].
We propose a novel, simple, and yet very effective algorithm that will stealthily and strategically manipulate the observations of compromised agents. The goal is to add small perturbations to these observations, causing the agents to perceive the same environment differently- a phenomenon we refer to as misalignment. For instance, imagine a simplified version of the Pursuit game [5], where two agents must cooperate to surround a moving object to receive a reward. If each agent receives conflicting information about the object’s position due to an adversarial attack, their ability to coordinate becomes severely compromised. Our crafted perturbations are designed specifically to induce such misalignment to the targeted agents. We refer to this attack as Align attack.
@@
The second main contribution of our paper is to propose an even more constrained adversarial attack: we consider an adversary with no access whatsoever (no observations, actions, or policies), possessing only the capability to inject small perturbations into each agent’s observation stream. The literature typically refers to this type of attack as Free attacks, which predominantly consist of injecting random noise. However, in our paper, we use the same idea of introducing misalignment in agents and make use of partial Hadamard matrices to generate better perturbations. As we explain later, misalignment can be induced using orthogonal matrices. We refer to this attack as Hadamard attack.
@@
Additionally, we propose a targeted attack that combines the two approaches: we use observations to select a subset of critical agents, and then use Hadamard matrices to craft adversarial perturbations. This allows us to benefit from the profiling capability of the Align attack and the fast generation of the Hadamard attack, thereby resulting in an efficient and lightweight attack.
Our contributions are illustrated in Figure 1. Most previous work considered attacks that rely on having access to some form of policy: the deployed policy (i.e., white-box attacks), a surrogate policy trained from scratch in the same environment, or a policy trained using imitation learning. These approaches depend on the transferability [19] of adversarial examples and tend to be very effective. However, as previously mentioned, they rely on strong assumptions about the attacker’s access, which reduces their plausibility. In contrast, we focus on threat models that assume weak access, which we believe are more realistic and yet remain underexplored in the literature. We argue that, in general, the fewer elements the attacker is assumed to access, the more plausible the attack becomes. Having access to observations only is more plausible than having access to both observations and actions, and both are more realistic than assuming access to the policy. Assuming no access at all is the most plausible scenario. We can summarize this intuition with the following relationship (The curve shown in Figure 1):
$\text{Plausibility}\approx\frac{1}{\text{Assumption strength}}$ (1)
It is important to note, however, that assuming no access does not mean the attack is trivial to implement. Like all the other threat models, we still assume the attacker is somehow present in the system and can inject perturbations. We only focus on what information the attacker relies on to generate those perturbations.
Finally, we demonstrate the effectiveness of our approach across three MARL benchmarks and 22 different tasks, covering fully-observable, partially-observable, and highly cooperative tasks. Our results show that our adversaries significantly undermine the performance of collaborative agents. Moreover, our attack applies to both value-based and policy-based algorithms. We also present a detailed analysis of the key factors that are particularly important for the performance and the computational cost of our attacks.
II Related works
Adversarial reinforcement learning can be classified according to several criteria: Is the system composed of a single agent or multiple agents? Does the attack occur during training or deployment? Are we considering a white-box or black-box scenario? What system components does the attacker have access to, and which does it target? Most existing work focuses on adversarial attacks against single-agent systems [18, 7, 8, 12, 22, 9, 28], often under white-box settings with access to states, actions, and rewards. In contrast, our work addresses adversarial attacks on deployed multi-agent systems in a black-box setting. We consider scenarios where the attacker initially has access only to the agents’ observations, and further extend our approach to cases where the adversary has no access at all. Table I provides a comprehensive comparison between our work and previous work on adversarial c-MARL.
TABLE I: Comparison of our work and previous papers. We indicate whether the work considers test-time vs training-time attacks and black-box vs white-box settings. We report the attacker’s accessible elements and targeted components.
@@ II Related works
Pham et al. (2023) [23]
Lin et al. (2020) [11]
Nisioti et al. (2021) [17]
Hu and Zhang (2022) [6]
Chen et al. (2024) [2]
Liu and Lai (2023) [13]
Zheng et al. (2023) [33]
@@
Training-time vs Test-time attacks. Training-time attacks, also referred to as data poisoning attacks [24], occur when an adversary is present during the training. The aim of these attacks is to manipulate the agent into learning a target policy crafted by the attacker. This can be accomplished through reward poisoning [13, 31] or environment poisoning [25], as well as by manipulating observations or actions [6, 2, 33]. However, it is not always feasible for the adversary to interfere with the training process. Multi-agent systems are arguably more vulnerable when deployed in the real world. Attacks occurring during deployment are referred to as test-time attacks, where the goal is to degrade the performance of already-trained policies. This is typically achieved by exploiting the known vulnerabilities of neural networks to adversarial inputs [7, 27], mainly through observation manipulation [23, 11], but also through action manipulation [17]. Our work focuses on test-time attacks that target observations. Much of prior work (see Table I) on test-time attacks assumes access to numerous elements at the same time, such as the policy network, observation, actions, rewards, and even the training environment—assumptions that are unrealistic in actual deployment scenarios. Therefore, we extend previous work by considering more practical scenarios, where the adversary has limited access, either only to the agents’ observations or no access at all.
White-box vs Black-box. In white-box scenarios, the attacker knows the learning algorithm and has access to the policy weights and its architecture [2, 6, 17, 11, 23, 7]. While these attacks tend to be the most effective, it is impractical for the attacker to have the complete knowledge of the deployed policies. Conversely, black box settings allow for more relaxed assumptions [7]. Most existing work relies on learning a surrogate policy to exploit the transferability of adversarial examples [19]. This surrogate policy can be learned by training the model from scratch [7], but doing so requires access to the training environment. Alternatively, imitation learning can be used to approximate the policy [28, 8], which would necessitate access to observation-action pairs or the ability to query the deployed policy. However, neither of these approaches is applicable in our scenario, as we lack access to the training environment and actions, and we cannot query the model.
Finally, as shown in Table I, we are the only work that investigates test-time attacks in black-box settings with minimal accessible information.
III Background
Consider a neural network $f$ parameterized by $\theta$, which takes an input $x\in\mathcal{X}$ and outputs $y$. The network is trained to minimize a loss function J. The goal of an adversarial attack is to add a perturbation $\delta$ to the original input $x$ in order to maximize the loss $\text{J}(x+\delta,y;\theta)$ with respect to $\delta$. Moreover, this perturbation must be small enough to avoid detection, which can be achieved by constraining the $L_{\infty}$ norm of the perturbation within an attack budget $\epsilon$. The perturbation $\delta$ can be determined by solving the following optimization problem:
@@ III Background
A wide range of attack methods has been proposed for generating adversarial perturbations. Fast Gradient Signed Methods (FGSM) [4] is a straightforward technique that generates perturbations as follows:
@@
FGSM is a single-step attack. A more powerful variant is the multistep attack known as Projected Gradient Descent (PGD) [10, 15]. This attack iteratively injects perturbations over $K$ steps with a small step size $\alpha$. Moreover, to ensure that the perturbed input remains within the valid input domain $\mathcal{X}$, the perturbed data is projected back into the domain after each step. PGD perturbs the inputs as follows:
@@
IV Methodology
IV-A Problem statement
We consider a cooperative multi-agent system during its deployment phase. Let $\mathcal{N}=\{1,\dots,n\}$ be the set of agents. We assume that an adversary has successfully compromised the entire set $\mathcal{N}$ or a subset of it. Although the adversary has compromised these agents, it does not possess direct control over their actions. Instead, for each compromised agent $i\in\mathcal{N}$, the adversary can inject adversarial perturbations before they are processed by the agent’s policy. The attacker must ensure that these perturbations remain small enough to avoid detection.
The main problem we seek to address can be outlined as follows:
Given the local observations of the agents, can an adversary undermine the performance of a deployed cooperative multi-agent system? And if these observations are inaccessible, can we still design effective noise-based adversarial attacks?
In the following paragraphs, we first explain the intuition behind our paper and then present our adversarial algorithms.
IV-B Our intuition
In cooperative tasks, the ability of agents to work together relies heavily on having aligned perceptions and beliefs of their environment. By aligned perceptions, we mean that agents observing the same object must perceive it consistently, which includes receiving identical information about its characteristics. For example, in the SMAC games [3], agents work together to defeat a set of enemy units in different battles. To succeed, agents must learn a range of important skills such as focus fire, where agents coordinate to jointly attack a single opponent to eliminate it quickly (see Figure 2). This requires the agents to agree on which enemy to jointly attack, typically selecting the weakest one (i.g, based on the health level). Thus, to properly use this skill, the team members should all have aligned perceptions of their targeted opponent, including its health and position. If this alignment is lacking, due to the adversary tampering with health values as in Figure 2, agents may fail to coordinate their attacks. This illustrates how misaligned perceptions can hinder agents’ ability to collaborate effectively and achieve shared goals, potentially pushing them to take opposing actions. Thus, if an adversary manages to properly manipulate agents’ perceptions, the collaboration is likely to be compromised.
@@ IV-B Our intuition
We do not necessarily assume that the agents must observe the same objects. The same intuition applies to partially observable environments where there is minimal to no overlap between agents’ observations. In these scenarios, our intuition is to induce misalignment in the common beliefs among the agents [30, 16]. Accordingly, our experimental section mainly focuses on partially observable tasks.
To summarize, the intuition behind our attack is to exploit this reliance on aligned perceptions and understandings among agents. We aim to perturb the observations of infected agents in a way that induces misalignment in their perceptions. This will create confusion and disagreements in how they understand their environment. To achieve this, we adopt a divide-and-conquer strategy by ensuring that each infected agent perceives the environment differently from the other compromised agents. The goal is to divide the compromised agents by creating discrepancies in their perceptions, severely undermining their ability to coordinate, which ultimately weakens the overall performance of the entire team.
In the following sections, we introduce two novel attacks. Sections IV-C and IV-D present the Align attack and its lightweight variant, respectively, while Section IV-E describes the Hadamard attack.
IV-C The Align attack
Figure 2: Illustration: a small adversarial perturbation (0.01,0,0.04)(0.01,0,-0.04) could prevent the agents from coordinating their actions effectively
The standard assumption in c-MARL systems is that collaborative agents typically have aligned perceptions, meaning they usually receive consistent information from their environments. Therefore, it is accepted to assume that the observations of the infected agents $\{o_{1},o_{2},\dots,o_{n}\}$ are aligned. Our goal is to manipulate each observation so that the new observations $\{o_{1}+\delta_{1},o_{2}+\delta_{2},\dots,o_{n}+\delta_{n}\}$ become misaligned, resulting in conflicting information among those agents. However, before we can do this, we must first be able to measure whether the observations are misaligned, as this will inform our attack.
When agents have aligned perceptions, their observations tend to correlate because they contain similar information or have comparable understandings of the environment. This means that the observations of one agent can be approximated based on the observations of other agents. On the other hand, when their observations are misaligned, they become less correlated and contain less overlapping content.
Formally, when agents’ perceptions are aligned, we can train the following neural network $f_{\theta}$:
@@ IV-C The Align attack
where $o_{-i}=\{o_{1},\dots,o_{i-1},o_{i+1},\dots o_{n}\}$ is the concatenated observations of all the agents except agent $i$. This network can be trained using the mean squared error:
@@
where $\mathbf{o}$ is the joint observation.
Once $f_{\theta}$ is properly trained, it can effectively reconstruct each agent’s observations using the observations from its peers, which will manifest as a low loss value. Conversely, when the network is fed with misaligned observations, we will notice high loss values. Therefore, a simple strategy for the adversary to induce misalignment to the agents is to add a small perturbation $\{\delta_{i}\}_{i\in\mathcal{N}}$ to the observations that will maximize the loss.
This strategy can be formally expressed as:

### v2

# Constrained Black-Box Attacks Against Cooperative Multi-Agent Reinforcement Learning

@@ Abstract
Collaborative multi-agent reinforcement learning has rapidly evolved, offering state-of-the-art algorithms for real-world applications, including sensitive domains. However, a key challenge to its widespread adoption is the lack of a thorough investigation into its vulnerabilities to adversarial attacks. Existing work predominantly focuses on training-time attacks or unrealistic scenarios, such as access to policy weights or the ability to train surrogate policies. In this paper, we investigate new vulnerabilities under more challenging and constrained conditions, assuming an adversary can only collect and perturb the observations of deployed agents. We also consider scenarios where the adversary has no access at all (no observations, actions, or weights). Our main approach is to generate perturbations that intentionally misalign how victim agents see their environment. Our approach is empirically validated on three benchmarks and 22 environments, demonstrating its effectiveness across diverse algorithms and environments. Furthermore, we show that our algorithm is sample-efficient, requiring only 1,000 samples compared to the millions needed by previous methods.
1 Introduction
@@
Collaborative multi-agent reinforcement learning (c-MARL) algorithms have demonstrated state-of-the-art performances in complex cooperative tasks Rashid et al. (2018); Yu et al. (2022), making them well-suited for solving real-world problems across various domains Lv et al. (2025); Park et al. (2024); Zhang et al. (2024). However, a critical prerequisite for the widespread adoption of c-MARL is a full understanding of its vulnerabilities to adversarial attacks Huang et al. (2017); Kos and Song (2017), particularly when deployed.
@@
While much of the literature on adversarial c-MARL focuses on training-time attacks Zheng et al. (2023); Liu and Lai (2023); Chen et al. (2024); Hu and Zhang (2022), we focus instead on test-time attacks, where the adversary is present during deployment. Prior work on test-time attacks Pham et al. (2023); Lin et al. (2020); Nisioti et al. (2021) has primarily considered white-box threat models, in which the adversary has access to the policy architecture and weights. This scenario is not always feasible. Moreover, such access can itself be considered a successful attack, as it typically involves proprietary knowledge with significant financial implications if leaked. In contrast, black-box threat models Huang et al. (2017) do not assume access to the policy’s weights or architecture; instead, they often involve learning a surrogate policy network either by training from scratch in the same environment Huang et al. (2017) or through imitation learning Wu et al. (2021); Inkawhich et al. (2020). The former requires access to the training environment, while the latter relies on collecting both observations and actions or the ability to query the model (see Figure 1). But what if the adversary cannot access such information to train a surrogate policy?
@@
In our work, we examine test-time attacks against c-MARL in black-box settings, but we push the standard assumptions a step further: we consider an adversary whose access is restricted to the observations (no access to the actions) of deployed agents, referred to as infected agents. The adversary does not have full control over these infected agents; instead, they can only add small perturbations to their observations.
Figure 1: Assumption Strength vs Plausibility.
Our first main contribution is to show that such an adversary can effectively sabotage a c-MARL system, causing significant damage with as few as 1,000 collected samples, compared to the millions required by previous methods Pham et al. (2023). Our idea is to generate small perturbations that cause agents to have inconsistent views of the same environment; we refer to this as misalignment. For instance, in a Pursuit game Gupta et al. (2017), where two agents must cooperate to surround a moving target, if an adversary causes each agent to observe a different target position, their ability to coordinate is severely compromised. We design our perturbations to induce this effect and refer to the resulting attack as the Align attack.
@@
Our second main contribution introduces an even more constrained attack. We consider an adversary with no access whatsoever (no observations, actions, or weights), having only the ability to perturb each agent’s observation. The literature refers to this setting as free attacks, which predominantly rely on random noise injection. In contrast, we apply the same misalignment principle and propose a novel attack using partial Hadamard matrices to generate structured perturbations. As discussed in Section 4.5, misalignment can be induced through orthogonal matrices. We refer to this attack as the Hadamard attack.
@@
Moreover, we propose a targeted attack that combines ideas from both methods by using observations to identify critical agents and Hadamard matrices to craft adversarial perturbations. This leverages the profiling capability of the Align attack and the fast generation of the Hadamard attack, resulting in an efficient and lightweight attack.
Our contributions are illustrated in Figure 1. Most prior work considers attacks that require access to some form of policy, real or surrogate, and tend to be effective. However, such methods rely on strong access assumptions, which limit their plausibility. In contrast, we study threat models with weak access, which are more challenging and remain underexplored. In general, fewer attacker assumptions lead to more plausible attacks. Figure 1 illustrates this inverse relationship between attack plausibility and threat model strength.
Finally, we conduct extensive experiments on three MARL benchmarks and 22 tasks, covering fully observable, partially observable, and highly cooperative settings, along with appropriate ablation studies.
2 Related works
Adversarial c-MARL can be classified according to several criteria: Does the attack occur during training or deployment? Are we considering a white-box or black-box scenario? What system components does the attacker have access to, and which does it target?
Table 1: Comparison with prior work.
@@ 2 Related works
Pham et al. (2023) Pham et al. (2023)
Lin et al. (2020) Lin et al. (2020)
Nisioti et al. (2021) Nisioti et al. (2021)
Chen et al. Chen et al. (2025)
Hu and Zhang (2022) Hu and Zhang (2022)
Chen et al. (2024) Chen et al. (2024)
Liu and Lai (2023) citeliu2023efficient
Zheng et al. (2023) Zheng et al. (2023)
@@
Training-time vs Test-time attacks. Training-time attacks, also called data poisoning attacks Rakhsha et al. (2020), occur when an adversary is present during the training and aim to manipulate the agent into a target policy. This can involve reward poisoning Liu and Lai (2023); Zhang et al. (2020), environment poisoning Rakhsha et al. (2021), or altering observations or actions Hu and Zhang (2022); Chen et al. (2024); Zheng et al. (2023). However, interfering with training is not always feasible. MARL systems are arguably more vulnerable when deployed. Test-time attacks occur during deployment and aim to degrade agent performance. This is achieved by exploiting the known vulnerabilities of neural networks to adversarial inputs Huang et al. (2017); Szegedy (2013), mainly through observation manipulation Pham et al. (2023); Lin et al. (2020), or action manipulation Nisioti et al. (2021).
White-box vs Black-box. In white-box scenarios, the attacker knows the learning algorithm and has access to the policy weights and its architecture Chen et al. (2024); Hu and Zhang (2022); Nisioti et al. (2021); Lin et al. (2020); Pham et al. (2023); Huang et al. (2017). While these attacks tend to be the most effective, it is impractical for the attacker to have the complete knowledge of the deployed policies. Conversely, black box settings allow for more relaxed assumptions Huang et al. (2017). Most existing work relies on learning a surrogate policy to exploit the transferability of adversarial examples Papernot et al. (2016). This surrogate policy can be learned by training the model from scratch Huang et al. (2017), but doing so requires access to the training environment. Alternatively, imitation learning can be used to approximate the policy Wu et al. (2021); Inkawhich et al. (2020), which would necessitate access to observation-action pairs or the ability to query the policy.
Table 1 provides a comprehensive comparison between our work and previous work. Prior work on test-time attacks often assumes white-box adversaries and access to multiple elements simultaneously, which is not always feasible during deployment. We instead focus on more practical scenarios: deployed c-MARL in a black-box setting with limited access.
3 Background
Consider a neural network $f$ parameterized by $\theta$, which takes an input $x\in\mathcal{X}$ and outputs $y$, and is trained to minimize a loss function J. An adversarial attack aims to add a perturbation $\delta$ to the input $x$ that maximizes the loss $\text{J}(x+\delta,y;\theta)$. Moreover, $\delta$ must remain small to avoid detection, which is enforced by constraining its $L_{\infty}$ norm with a budget $\epsilon$. This problem is formulated as:
@@ 3 Background
Many methods have been proposed to find $\delta$. Fast Gradient Signed Methods (FGSM) Goodfellow et al. (2014) is a simple technique that computes the perturbations as:
@@
FGSM is a single-step attack. A stronger multistep variant is Projected Gradient Descent (PGD) Kurakin et al. (2018); Madry et al. (2018) which applies iterative perturbations over $K$ steps with a small step size $\alpha$. After each step, the perturbed input is projected back into the valid domain $\mathcal{X}$ :
@@
Using FGSM or PGD against a c-MARL system requires access to the agents’ policy or value networks in order to compute the loss in Equation 1. Since we consider attacks in a black-box setting, the main challenge our paper tries to solve is how to circumvent the need for direct access to the policy and still be able to use powerful methods like PGD.
4 Methodology
4.1 Problem statement
We consider a cooperative multi-agent system during its deployment phase. Let $\mathcal{N}=\{1,\dots,n\}$ be the set of agents. We denote by $o_{i}$ the observation of agent $i$, which is of dimension $d$, $o_{-i}$ is the concatenated observations of all the agents except agent $i$, and $\mathbf{o}$ is the joint observation. The adversary cannot directly control the agents’ actions but can inject small perturbations into their observations before the policy networks process them. This assumption is standard in adversarial c-MARL (see Section 2).
4.2 Intuition
Figure 2: Coordination is prevented by flipping the lowest-health enemy as seen by each agent.
In cooperative tasks, effective collaboration depends on agents having aligned perceptions and beliefs about their environment. By aligned perceptions, we mean that agents observing the same object receive consistent information about its attributes. For instance, in the SMAC benchmark Ellis et al. (2024), agents cooperate to defeat enemy units, which requires skills such as focus fire, where agents jointly attack a single opponent to eliminate it quickly . This typically involves selecting the weakest enemy based on health level. Effective coordination therefore requires agreement on the target’s health and position. Adversarial manipulation of these perceptions, as shown in Figure 2, can disrupt coordination.
@@ 4.2 Intuition
We do not necessarily assume that the agents must observe the same objects. The same intuition applies to partially observable environments where there is minimal to no overlap between agents’ observations. In this case, our goal is to induce misalignment in the common beliefs among the agents Zhang et al. (2022); Moreno et al. (2021). Accordingly, our experiments mainly focus on partially observable tasks.
Overall, our attack exploits agents’ dependence on aligned perceptions by manipulating their observations. Our idea is similar to a divide-and-conquer strategy: we make each agent perceive the environment differently, which undermines coordination and ultimately degrades team performance.
4.3 The Align attack
Cooperative MARL typically assumes that agents have aligned perceptions. Under this assumption, agents’ observations $\{o_{1},\dots,o_{n}\}$ are initially aligned. Our goal is to perturb each observation so that the modified observations $\{o_{1}+\delta_{1},\dots,o_{n}+\delta_{n}\}$ become misaligned. Therefore, we first need a quantitative measure of misalignment.
Aligned observations are correlated because they encode similar information and beliefs about the environment. As a result, one agent’s observation $o_{i}$ can be approximated from the observations of other agents $o_{-i}$. In contrast, misaligned observations exhibit weaker correlation, making such an approximation difficult.
Formally, aligned observations allow us to train the following neural network $f_{\theta}$ using the mean squared error $\text{J}(\mathbf{o};\theta)$ :
@@ 4.3 The Align attack
Once $f_{\theta}$ is properly trained, it can be used to measure misalignment, as high loss values are associated with misaligned observations. Therefore, our strategy to induce misalignment is to find small perturbations $\{\delta_{i}\}_{i\in\mathcal{N}}$ that will maximize the loss of the trained network $f_{\theta}$. This is formally expressed as: